40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 23 15:16:52 2017, gautam, Update, Optical Levers, ETMY Oplev HeNe needs to be replaced ETMY_Oplev.pdf
    Reply  Mon Jan 23 17:24:26 2017, Steve, Update, Optical Levers, ETMY Oplev HeNe is running hot ETMY_oplev__lasers.png
       Reply  Tue Jan 24 17:52:15 2017, Craig, Update, Optical Levers, ETMY Oplev HeNe is replaced 6x
          Reply  Wed Mar 1 08:28:04 2017, Steve, Update, Optical Levers, oplev laser RIN RIN_1103P.pngRIN_HeNe.png
             Reply  Thu Mar 2 13:59:04 2017, Steve, Update, Optical Levers, oplev laser RIN 1103P@2d.png
                Reply  Thu Mar 2 17:58:45 2017, rana, Update, Optical Levers, oplev laser RIN 
                   Reply  Fri Mar 3 17:24:21 2017, Steve, Update, Optical Levers, oplev laser RIN 3march17.pdf
                      Reply  Wed Mar 8 15:28:37 2017, Steve, Update, Optical Levers, oplev laser RIN RIN.jpgRIN_1103P_rotated.pdf
                      Reply  Mon Apr 10 09:49:32 2017, Steve, Update, Optical Levers, oplev laser RIN test planning 
                         Reply  Mon Apr 10 14:21:57 2017, rana, Update, Optical Levers, oplev laser RIN test planning 
       Reply  Tue May 30 13:36:58 2017, Steve, Update, Optical Levers, ETMY Oplev HeNe is replaced oplevETMY120d.png
          Reply  Tue Jan 15 18:16:00 2019, gautam, Update, Optical Levers, ETMY Oplev HeNe needs replacement OLRIN.pdfOLsums.png
             Reply  Fri Jan 18 12:52:07 2019, gautam, Update, Optical Levers, BS/PRM Oplev HeNe replaced OLRIN_20190118.pdf
Message ID: 12873     Entry time: Wed Mar 8 15:28:37 2017     In reply to: 12866
Author: Steve 
Type: Update 
Category: Optical Levers 
Subject: oplev laser RIN 

Gautam and Steve,

Quote:

Corrected oplev laser RIN plot at day 3

RXA:

  1. to measure RIN, the lever arm should be really short, not long.
  2. the beam should be 3x smaller than the active area of the diode
  3. the specular beam should be dumped on a razor dump.
  4. we need to make a summary page for HeNe laser testing so that we can see 24 hour specgrams of these things for ~3-4 lasers at the same time.
  5. We should add specgram stuff for the existing HeNe SUM channels on the active OLs.

GV: The channel the PD Steve is using is hooked up to C1:ALS-FC_X_F_IN. As I found out today, there can be considerable RF pickup between the C1:ALS-FC_X_F_IN and C1:ALS-FC_Y_F_IN channels, which share a common 4-pin LEMO cable - this is because the rise time of the square wave output of the Wenzel dividers is <1us, so suitability of this particular channel for the RIN measurement set up has to be reconsidered. Perhaps we can use one of the six spare PEM channels over at 1X6. 

We did the following:

1, switched data channel  from  C1:ALS-FC_X_F_IN to C1:PEM-MIC_1_OUT_DQ   Actual connection at 1X7 rack, input C17

    Tested channel with 1Hz, 100 mV sine wave through DV

2, placed BS into the beam path so the reflected value on the PDA100A 0.1mW,  beam od ~1mm, beam path lenght 11 cm, gain 20dB 3.7Vdc

    The full output of this 1103P 2.8 mW was saturating the PDA100A

Summery :finding it to be too good to be this good

Attachment 1: RIN.jpg  211 kB  | Hide | Hide all
RIN.jpg
Attachment 2: RIN_1103P_rotated.pdf  37 kB  | Hide | Hide all
RIN_1103P_rotated.pdf
ELOG V3.1.3-